# RBI-ThPhys-2024-15

### v1

@@ VII Interference effects on differential distributions
We use the scans shown in Sect. V to define several interesting benchmark (BM) points. We select the BM points such that they exhibit interesting features that we will describe below. Additionally, we ensure that all points are accessible to experimental searches at the LHC, by requiring that the $gg\rightarrow H\rightarrow hh$ ($\mathrm{S_{H}}$-only) cross-section $\sigma_{\mathrm{S_{H}}}$, scaled by the k-factors as described in Sect. VI, is close to the experimental limits presented by the ATLAS Collaboration in Ref. ATLAS:2021tyg . To achieve this we require $\sigma_{\mathrm{S_{H}}}>\sigma_{\mathrm{ATLAS}}\times\sqrt{\frac{140}{L[\mathrm{fb}^{-1}]}}$, where $\sigma_{\mathrm{ATLAS}}$ is the expected limits from Ref. ATLAS:2021tyg , and $L$ is a target luminosity [^15] We chose to use the ATLAS limits as no Run-2 CMS combination was published at the time. Since then, the CMS Collaboration has presented a combined result in Ref. CMS:2024phk . We have checked whether taking the best expected limit from either CMS or ATLAS has any influence on the defined BM points, and confirmed that the same BMs are selected in this case.. We consider two values of $L$ equal to 400 $\mathrm{fb}^{-1}$ and 3000 $\mathrm{fb}^{-1}$, which approximately correspond to the expected luminosity recorded by each of the ATLAS/CMS detectors at the end of Run-3 (assuming Run-2+Run-3 are combined) and at the end of the LHC program, respectively.
Tab. 4 summarizes the BM points, which will be described in more detail in the following sections. All points in the table satisfy the $\sigma_{\mathrm{S_{H}}}$ requirement for $L=3000~{}\mathrm{fb}^{-1}$, and we additionally indicate in the second to last column if the requirement is also satisfied for $L=400~{}\mathrm{fb}^{-1}$. In the following sections, we will state the parameters of the singlet model for the chosen benchmark scenarios and show distributions produced from MC events generated at LO in QCD.
@@
Table 4: Overview of the investigated benchmark points. The values of the $\sin\alpha$, $\tan\beta$, and $m_{H}$ model parameters are shown. The values of $\Gamma_{H}$, $\kappa_{\lambda_{hhh}}$, $\sigma$, and $\sigma_{\mathrm{S_{H}}}$, which are fixed by the model parameters, are also displayed. The cross-sections are computed at LO and scaled by NNLO k-factors as described in Sect. VI. The second to last column indicates whether the points are expected to be accessible at the LHC by the end of Run-3. The last column gives a summary of the feature exhibited by the benchmark point.
@@ VII.1 BM1
The $m_{hh}$ distributions for this BM with and without detector smearing are displayed in Fig. 18. Smearing in this and all subsequent figures is performed according to the prescription discussed in Sect. IV. The figure compares the distributions for four different treatments of the di-Higgs modelling. The full di-Higgs distribution including both resonant and non-resonant diagrams and interference terms is shown in blue, the distribution neglecting the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms in shown in red (refereed to as “no-interference” in the following), the $\mathrm{S_{H}}$-only distribution is shown in green, and the distribution obtained by fixing the non-resonant di-Higgs spectrum to the SM expectation and summing with the $\mathrm{S_{H}}$ distribution is shown in magenta (referred to as “SM+$\mathrm{S_{H}}$”).
@@
The interference tends increase (decrease) the broader non-resonant-like continuum above (below) the resonant peak at $m_{H}\sim 620\,{\,\rm GeV}$. After detector smearing, the overall distribution exhibits an interesting double-peak structure. The interference also causes a shift in the position of the mass peak to slightly lower values, however, after smearing the shift in the peak position is not very noticeable. It is clear from comparing to the $\mathrm{S_{H}}$-only spectrum that neglecting the non-resonant terms altogether results in a failure to describe the double-peak structure, and it also results in underestimation of the number of events close to the peak of around 35%. Including the non-resonant terms while neglecting only the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms clearly brings an improvement in the overall description of the spectrum. However, it still results in an underestimate of the height of the $\mathrm{S_{H}}$ peak by about 15%, and an under(over)-estimate of the non-resonant continuum of $\approx$ 15–30% (15–40%) below (above) the resonant peak. This clearly motivates the proper consideration of the interference effects when performing resonant searches. We also do not observe a significant difference between no-interference and SM+$\mathrm{S_{H}}$ distributions. This is because only small values of $\sin\alpha$ are still allowed by the experimental constraints and this results in $\kappa_{\lambda_{hhh}}$ and $\kappa_{q}^{h}$ values that are typically close to unity. As will be shown in the subsequent sections, this feature is common to all BM points.
@@
Figure 18: The di-Higgs mass distributions for BM1 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@
We also show the distributions for several other variables in Figs. 19-28. The variables displayed are: $p_{\mathrm{T}}^{hh}$, the $p_{\mathrm{T}}$ of the leading and sub-leading $h$, $p_{\mathrm{T}}^{h_{1}}$ and $p_{\mathrm{T}}^{h_{2}}$, the $p_{\mathrm{T}}$ of the leading and softest $b$-jets, $p_{\mathrm{T}}^{b_{1}}$ and $p_{\mathrm{T}}^{b_{4}}$, the $\mathrm{H}_{\mathrm{T}}$ of the four $b$-jets, the $\eta$ of the di-Higgs system, $\eta^{hh}$, the separation in $\eta$ of the two $h$, $\Delta\eta(h_{1},h_{2})$, the separation of the $\phi$ between the two $h$, $\Delta\phi(h_{1},h_{2})$, and the $\Delta R$ (where $\Delta R^{2}=\Delta\phi^{2}+\Delta\eta^{2}$) between the two $h$, $\Delta R(h_{1},h_{2})$. In general, the effect of the interference on the shapes of these variables is less significant than was observed for $m_{hh}$. The $p_{\mathrm{T}}^{hh}$ and $\eta^{hh}$ distributions in particular do not show any noticeable differences in shapes when neglecting the interference terms. There are more noticeable differences in the other distributions, however, the regions which have largest differences tend to occur in the tails and are therefore probably do not have a very large impact on the experimental searches. The $\mathrm{S_{H}}$-only $p_{\mathrm{T}}^{h_{1}}$, $p_{\mathrm{T}}^{h_{2}}$, $p_{\mathrm{T}}^{b_{1}}$, $p_{\mathrm{T}}^{b_{4}}$, and $\mathrm{H}_{\mathrm{T}}$ distributions on the other hand do show very obvious differences compared to the full spectrum which further discourages neglecting the non-resonant contribution when performing experimental searches.
@@
Figure 19: The di-Higgs $p_{\mathrm{T}}$ distributions for BM1 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@
Figure 20: The leading $h$ $p_{\mathrm{T}}$ for BM1 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@
Figure 21: The sub-leading $h$ $p_{\mathrm{T}}$ before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@
Figure 22: The $p_{\mathrm{T}}$ of the leading $b$-jet before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta. No $p_{\mathrm{T}}$ or $\mathrm{H}_{\mathrm{T}}$ cuts have been applied for the smeared distributions in this case.
@@
Figure 23: The $p_{\mathrm{T}}$ of the softest $b$-jet before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta. No $p_{\mathrm{T}}$ or $\mathrm{H}_{\mathrm{T}}$ cuts have been applied for the smeared distributions in this case.
@@
Figure 24: The $\mathrm{H}_{\mathrm{T}}$ of the $b$-jets before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta. No $p_{\mathrm{T}}$ or $\mathrm{H}_{\mathrm{T}}$ cuts have been applied for the smeared distributions in this case.
@@
Figure 25: The di-Higgs $\eta$ distributions for BM1 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@
Figure 26: The $\Delta\eta$ between the two $h$ for BM1 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@
Figure 27: The $\Delta\phi$ between the two $h$ before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.
@@
Figure 28: The $\Delta R$ between the two $h$ before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $h$ s-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the superposition of $\mathrm{S_{H}}$ with the SM di-Higgs spectrum in magenta.

### v2

@@ VII Interference effects on differential distributions
We use the scans shown in Section V to define several interesting BM points. We select the BM points such that they exhibit interesting features that we will describe below. Additionally, we ensure that all points are accessible to experimental searches at the LHC, by requiring that the $gg\rightarrow H\rightarrow hh$ ($\mathrm{S_{H}}$-only) cross-section $\sigma_{\mathrm{S_{H}}}$, scaled by the K-factors as described in Section VI, is close to the experimental limits presented by the ATLAS Collaboration in Ref. ATLAS:2021tyg . To achieve this we require $\sigma_{\mathrm{S_{H}}}>\sigma_{\mathrm{ATLAS}}\times\sqrt{\frac{140}{L[\mathrm{fb}^{-1}]}}$, where $\sigma_{\mathrm{ATLAS}}$ is the expected limits from Ref. ATLAS:2021tyg , and $L$ is a target luminosity [^16] We chose to use the ATLAS limits as no Run-2 CMS combination was published at the time. Since then, the CMS Collaboration has presented a combined result in Ref. CMS:2024phk . We have checked whether taking the best expected limit from either CMS or ATLAS has any influence on the defined BM points, and confirmed that the same BMs are selected in this case.. We consider two values of $L$ equal to 400 $\mathrm{fb}^{-1}$ and 3000 $\mathrm{fb}^{-1}$, which approximately correspond to the expected luminosity recorded by each of the ATLAS/CMS detectors at the end of Run-3 (assuming Run-2+Run-3 are combined) and at the end of the LHC program, respectively.
Table 4 summarises the BM points, which will be described in more detail in the following sections. All points in the table satisfy the $\sigma_{\mathrm{S_{H}}}$ requirement for $L=3000~{}\mathrm{fb}^{-1}$, and we additionally indicate in the second to last column if the requirement is also satisfied for $L=400~{}\mathrm{fb}^{-1}$. In the following sections, we will state the parameters of the singlet model for the chosen benchmark scenarios and show distributions produced from MC events generated at LO in QCD. The K-factor scaling described in Section VI is also applied. We remind the reader than the shapes of the $m_{hh}$ distributions generated at LO were shown to agree very well with the NLO distributions, and thus we deem the LO samples to be sufficient for these comparisons. We note that the $p_{\mathrm{T}}^{hh}$ spectrum at LO does not accurately match the NLO predictions, therefore, any analysis relying on these or similar kinematic observables should be generated at NLO, as argued above.
@@
BM9 0.16 1.0 800 9.8 0.96 38.9 3.6 Max $\frac{\Gamma_{H}}{m_{H}}$
Table 4: Overview of the investigated benchmark points. The values of the $\sin\alpha$, $\tan\beta$, and $m_{H}$ model parameters are shown. The values of $\Gamma_{H}$, $\kappa_{\lambda_{hhh}}$, $\sigma$, and $\sigma_{\mathrm{S_{H}}}$, which are fixed by the model parameters, are also displayed. The cross-sections are computed at LO and scaled by NNLO K-factors as described in Section VI. The second to last column indicates whether the points are expected to be accessible at the LHC by the end of Run-3. The last column gives a summary of the feature exhibited by the benchmark point.
@@ VII.1 BM1
The $m_{hh}$ distributions for this BM with and without detector smearing are displayed in Figure 18. Smearing in this and all subsequent figures is performed according to the prescription discussed in Section IV. The figure compares the distributions for four different treatments of the di-Higgs modelling. The full di-Higgs distribution including both resonant and non-resonant diagrams and interference terms is shown in blue, the distribution neglecting the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms in shown in red (referred to as “no-interference” in the following), the $\mathrm{S_{H}}$-only distribution is shown in green, and the distribution obtained by fixing the non-resonant di-Higgs spectrum to the SM expectation and summing with the $\mathrm{S_{H}}$ distribution is shown in magenta (referred to as “SM+$\mathrm{S_{H}}$”).
@@
The interference tends increase (decrease) the broader non-resonant-like continuum above (below) the resonant peak at $m_{H}\sim 620\,{\,\rm GeV}$. After detector smearing, the overall distribution exhibits an interesting double-peak structure. The interference also causes a shift in the position of the mass peak to slightly lower values, however, after smearing the shift in the peak position is not very noticeable. It is clear from comparing to the $\mathrm{S_{H}}$-only spectrum that neglecting the non-resonant terms altogether results in a failure to describe the double-peak structure, and it also results in an underestimation of the number of events close to the peak of around 35%. Including the non-resonant terms while neglecting only the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms clearly brings an improvement in the overall description of the spectrum. However, it still results in an underestimate of the height of the $\mathrm{S_{H}}$ peak by about 15%, and an under(over)-estimate of the non-resonant continuum of $\approx$ 15–30% (15–40%) below (above) the resonant peak. This clearly motivates the proper consideration of the interference effects when performing resonant searches. We also do not observe a significant difference between no-interference and SM+$\mathrm{S_{H}}$ distributions. This is because only small values of $\sin\alpha$ are still allowed by the experimental constraints and this results in $\kappa_{\lambda_{hhh}}$ and $\kappa_{q}^{h}$ values that are typically close to unity at LO. As will be shown in the subsequent sections, this feature is common to all BM points.
@@
Figure 18: The di-Higgs mass distributions for BM1 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@
We also show the distributions for several other variables in Figures 19-28. The variables displayed are: $p_{\mathrm{T}}^{hh}$, the $p_{\mathrm{T}}$ of the leading and sub-leading $h$, $p_{\mathrm{T}}^{h_{1}}$ and $p_{\mathrm{T}}^{h_{2}}$, the $p_{\mathrm{T}}$ of the leading and softest $b$-jets, $p_{\mathrm{T}}^{b_{1}}$ and $p_{\mathrm{T}}^{b_{4}}$, the $\mathrm{H}_{\mathrm{T}}$ of the four $b$-jets, the $\eta$ of the di-Higgs system, $\eta^{hh}$, the separation in $\eta$ of the two $h$, $\Delta\eta(h_{1},h_{2})$, the separation of the $\phi$ between the two $h$, $\Delta\phi(h_{1},h_{2})$, and the $\Delta R$ (where $\Delta R^{2}=\Delta\phi^{2}+\Delta\eta^{2}$) between the two $h$, $\Delta R(h_{1},h_{2})$. In general, the effect of the interference on the shapes of these variables is less significant than was observed for $m_{hh}$. The $p_{\mathrm{T}}^{hh}$ and $\eta^{hh}$ distributions in particular do not show any noticeable differences in shapes when neglecting the interference terms. There are more noticeable differences in the other distributions, however, the regions which have largest differences tend to occur in the tails and therefore probably do not have a very large impact on the experimental searches. The $\mathrm{S_{H}}$-only $p_{\mathrm{T}}^{h_{1}}$, $p_{\mathrm{T}}^{h_{2}}$, $p_{\mathrm{T}}^{b_{1}}$, $p_{\mathrm{T}}^{b_{4}}$, and $\mathrm{H}_{\mathrm{T}}$ distributions on the other hand do show very obvious differences compared to the full spectrum which further discourages neglecting the non-resonant contribution when performing experimental searches.
@@
Figure 19: The di-Higgs $p_{\mathrm{T}}$ distributions for BM1 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@
Figure 20: The leading $h$ $p_{\mathrm{T}}$ for BM1 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@
Figure 21: The sub-leading $h$ $p_{\mathrm{T}}$ before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@
Figure 22: The $p_{\mathrm{T}}$ of the leading $b$-jet before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta. No $p_{\mathrm{T}}$ or $\mathrm{H}_{\mathrm{T}}$ cuts have been applied for the smeared distributions in this case.
@@
Figure 23: The $p_{\mathrm{T}}$ of the softest $b$-jet before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta. No $p_{\mathrm{T}}$ or $\mathrm{H}_{\mathrm{T}}$ cuts have been applied for the smeared distributions in this case.
@@
Figure 24: The $\mathrm{H}_{\mathrm{T}}$ of the $b$-jets before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta. No $p_{\mathrm{T}}$ or $\mathrm{H}_{\mathrm{T}}$ cuts have been applied for the smeared distributions in this case.
@@
Figure 25: The di-Higgs $\eta$ distributions for BM1 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@
Figure 26: The $\Delta\eta$ between the two $h$ for BM1 before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@
Figure 27: The $\Delta\phi$ between the two $h$ before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.
@@
Figure 28: The $\Delta R$ between the two $h$ before (left) and after (right) experimental smearing. The blue lines show the full di-Higgs spectrum including all interference terms. This is compared to the spectrum excluding the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms shown in red, the $H$ $s$-channel spectrum ($\mathrm{S_{H}}$) shown in green, and the incoherent sum of $\mathrm{S_{H}}$ and the SM di-Higgs spectrum in magenta.